OpenStudy (anonymous):

A curve with the equation y= -(x+p)(x+q) cuts the x-axis at the points A(-2,0) and B(6,0). Write down the values of p and q.

OpenStudy (unklerhaukus):

the curve cuts the axis when y=0

OpenStudy (anonymous):

So sub A(-2,0) into the equation? I got stuck because there were too many unknowns :/

OpenStudy (mertsj):

Write a system of two equations. For the first equation, replace x with -2 and y with 0. For the second equation replace x with 6 and y with 0. Now you have two equations with variables p and q. Solve the system.

OpenStudy (anonymous):

I got stuck again. Ended up with q+p=-4 .

hartnn (hartnn):

p+q=-4 is correct, now substitute back in pq+36+6(p+q)=0 to get pq=-12 can u find p and q knowing p+q=-4 and pq= -12 ???

OpenStudy (anonymous):

yeah I got it (: Thank you!

hartnn (hartnn):

welcome :)